# SincVAE: a New Approach to Improve Anomaly Detection on EEG Data Using SincNet and Variational Autoencoder

### v1

@@ 5 Results
In this section, the results of the seizure detection experiments conducted on both the Bonn and CHB-MIT datasets are presented. For each dataset, the models were trained using the configurations identified during their respective model selection stages. During each training phase, 20 % of the training data was randomly sampled and used as the validation set.
It is reminded that for each results, the model without the SincNet layer is referred to as VAE; whereas, the model incorporating a SincNet layer on top the encoder network, which is the proposal of this work, is referred to as SincVAE. It is crucial to note that the primary focus of these experiments is to integrate SincNet into the VAE framework and evaluate its effectiveness in the tackled problem, rather than achieving new state-of-the-art results on the datasets used.
@@ 5.1 The Bonn Dataset
The results of the seizure detection for both declared comparisons, i.e. Set A vs Set E and Set B vs Set E, will be discussed. In both scenarios, the test set includes 2760 samples, which exhibits a significant imbalance with 460 samples labeled as non-seizure and 2300 labeled as seizure. This disparity was taken in account during the evaluations.
Figure 3 shows the MSE distributions for the seizure data (top-left) and non-seizure data from (top-right) in the test sets, as well as for the validation data (bottom), comparing both the VAE and SincVAE models. The summary statistics for these MSE distributions are detailed in Table 6.
| Method | Seizure Test MSE | Non-Seizure Test MSE | Validation Set MSE | ||||||
| Mean | Min | Max | Mean | Min | Max | Mean | Min | Max | |
| SincVAE | |||||||||
| VAE | |||||||||
It is observed that, on non-seizure test data, the SincVAE model exhibits a mean MSE of , while the VAE shows a slightly lower mean MSE of . Both results align with the statistics from the validation set.
Conversely, on the seizure test data, SincVAE shows a mean MSE of , indicating significant variability among cases, whereas VAE presents a lower mean MSE of . SincVAE’s larger discrepancy between its seizure and non-seizure MSE values could indicate a better distinction between the seizure and non-seizure conditions, which could ease the threshold-based classification. In contrast, the narrower MSE range of VAE could make this differentiation more challenging. Similar conclusions can be drawn on Set B (see Figure 4 and Table 7).
| Method | Seizure Test MSE | Non-Seizure Test MSE | Validation Set MSE | ||||||
| Mean | Min | Max | Mean | Min | Max | Mean | Min | Max | |
| SincVAE | |||||||||
| VAE | |||||||||
To assess the classification performance of the models, a classification threshold must be determined. This threshold is typically based on the MSE values derived from the trained model on the validation or training set, as suggested by various studies [97, 98, 99, 100], and tailored to meet specific user requirements. For this study, aiming to reduce the number of false positives, the classification threshold was selected using the following criteria:
@@
Figure 5 shows the confusion matrices for both SincVAE and VAE, evaluated under the two thresholds and , on the test data for the Set A vs Set E case.
@@
Both models demonstrate to correctly identificate non-seizure instances, as reflected by high values of true negatives across both thresholds. VAE slightly outperforms SincVAE in this aspect, primarily due to its tendency to generate fewer false positives. Notably, SincVAE shows fewer false negatives compared to the VAE model. This suggests that SincVAE may be more precise in predicting seizures, aligning with the observations made in 6. Table 8 shows F1, precision and recall metrics [61] related to this set of experiments.
| Method | Threshold set to | Threshold set to | ||||
@@
| SincVAE | 0.998 | 1 | 0.996 | 0.994 | 0.989 | 1 |
| VAE | 0.956 | 1 | 0.917 | 0.992 | 0.992 | 0.992 |
The performance of SincVAE appears to be superior to that of VAE on this experimental case: SincVAE demonstrates higher recall, crucial for ensuring no seizure goes undetected, and F1 scores, indicating it is the more reliable for the seizure detection across the chosen thresholds.
@@
Regarding the case Set B vs Set E, Table 9 shows the same classification metrics, while Figure 6 shows the confusion matrices. In this case, SincVAE consistently maintains high performance across both thresholds, demonstrating a lower number of both false positives and false negatives. This indicates a robust ability of SincVAE to accurately differentiate between seizure and non-seizure events.
| Method | Threshold set to | Threshold set to | ||||
@@
| SincVAE | 0.967 | 1 | 0.936 | 0.989 | 0.985 | 0.994 |
| VAE | 0.948 | 1 | 0.902 | 0.980 | 0.980 | 0.979 |
The SincVAE model consistently outperforms in terms of F1 score and recall across both scenarios and thresholds, indicating robust predictive capabilities, particularly in minimizing missed seizures, which could be a critical aspect for clinical applications, due to its impact on enhancing patient safety and treatment efficacy.
Figure 8 shows the EEG signal reconstructions performed by both the VAE and SincVAE models on seizure samples from Set E. Notably, the VAE model exhibits superior reconstruction fidelity for these seizure signals compared to SincVAE, which results in a higher number of false negatives. Consistent with the classification results, this difference in reconstruction quality suggests varying performance between the two models in processing seizure-related EEG data. The SincVAE’s difficulty in reconstructing seizure signals, suggests its ability to distinguish between normal and seizure states, thereby establishing it as a more reliable solution for seizure detection.
@@ 5.2 The CHB-MIT Dataset
Based on the findings from the experiments with the Bonn dataset, the analyses on the CHB-MIT dataset will be presented in terms of F1 score, precision, and recall performances of both the SincVAE and VAE models for each subject in the CHB-MIT dataset. The decision threshold was chosen to be unique and equal to the 95th percentile of the validation set’s MSE for all subjects Results are shown in Table 10.
| Subject | F1 | Precision | Recall | |||
| SincVAE | VAE | SincVAE | VAE | SincVAE | VAE | |
| 1 | 0.919 | 0.918 | 0.900 | 0.909 | 0.939 | 0.928 |
| 2 | 0.756 | 0.708 | 0.651 | 0.627 | 0.901 | 0.813 |
| 3 | 0.950 | 0.943 | 0.927 | 0.922 | 0.975 | 0.965 |
| 4 | 0.718 | 0.645 | 0.644 | 0.617 | 0.812 | 0.675 |
| 5 | 0.956 | 0.934 | 0.946 | 0.942 | 0.966 | 0.927 |
| 6 | 0.108 | 0.069 | 0.208 | 0.167 | 0.072 | 0.043 |
| 7 | 0.962 | 0.931 | 0.981 | 0.980 | 0.945 | 0.886 |
| 8 | 0.768 | 0.799 | 0.808 | 0.847 | 0.731 | 0.756 |
| 9 | 0.849 | 0.923 | 0.747 | 0.896 | 0.984 | 0.952 |
| 10 | 0.959 | 0.878 | 0.998 | 0.994 | 0.924 | 0.785 |
| 11 | 0.975 | 0.968 | 0.986 | 0.987 | 0.964 | 0.949 |
| 12 | 0.175 | 0.161 | 0.438 | 0.304 | 0.109 | 0.109 |
| 13 | 0.157 | 0.165 | 0.704 | 0.714 | 0.088 | 0.093 |
| 14 | 0.308 | 0.308 | 0.509 | 0.509 | 0.220 | 0.220 |
| 15 | 0.660 | 0.606 | 0.926 | 0.922 | 0.513 | 0.452 |
| 16 | 0.055 | 0.053 | 0.042 | 0.040 | 0.079 | 0.079 |
| 17 | 0.207 | 0.096 | 0.944 | 0.833 | 0.116 | 0.051 |
| 18 | 0.853 | 0.854 | 0.946 | 0.950 | 0.776 | 0.776 |
| 19 | 0.859 | 0.881 | 0.779 | 0.828 | 0.958 | 0.941 |
| 20 | 0.698 | 0.670 | 0.872 | 0.884 | 0.582 | 0.539 |
| 21 | 0.266 | 0.091 | 0.547 | 0.500 | 0.176 | 0.050 |
| 22 | 0.927 | 0.902 | 0.953 | 0.951 | 0.902 | 0.858 |
| 23 | 0.817 | 0.854 | 0.895 | 0.946 | 0.752 | 0.779 |
| 24 | 0.670 | 0.683 | 0.708 | 0.752 | 0.635 | 0.626 |
@@
Observing the F1 scores, SincVAE performs comparably or slightly better than VAE for the majority of the subjects in the CHB-MIT dataset. Specifically, for subjects 1, 2, 3, 4, 5, 6, 7, 10, 11, 12, 14, 15, 16, 17, 20, 21, and 22, SincVAE either matches or exceeds the performance of VAE. Notably, there are cases where VAE has an higher F1 score than SincVAE, specifically for subjects 8, 9, 13, 18, 19, 23, and 24. These variations suggest that VAE might be better suited for certain cases.
@@
It’s noteworthy that the F1 scores for some subjects, such as 6 and 13, are low, which may be influenced by factors such as the inherent quality of the EEG recordings, the need of specific data preprocessing or the need of more sophisticated model architectures. However, the main aim of this study is to explore the effectiveness of integrating a SincNet layer within a VAE framework, rather than establishing new state-of-the-art results on the CHB-MIT dataset. Thus, discussions on specific architectural enhancements to improve EEG data quality are outside the scope of this work, which focuses instead on assessing the added value of the SincNet layer within the VAE architecture.
@@
As stated above, the seizure tracks in the CHB-MIT dataset are included of preictal, ictal, and postictal phases. Figure 9 shows the MSE values obtained for each second of track number 19 from subject 9, a track identified by the dataset’s authors as including a seizure event from seconds 5299 to 5361. These specific seconds are visually demarcated with two dashed vertical blue lines. The horizontal red line on the plot indicates the decision threshold, set at the 95th percentile of the MSE values obtained from the validation set. Blue dots on the plot represent EEG recordings that have been classified as non-seizure, while red dots indicate those classified as seizure.
It can be noted that several recordings are classified as anomalous by both of the models either in the preictal and in postictal phase. Figure 10 shows the percentage of anomalous point detected by SincVAE and VAE in the preictal, ictal and postictal phases.
Both models exhibit detect seizures during the ictal phase with a high rate, and demonstrate effectiveness in recognizing the interictal phase with a low rate of false positives. It is interesting to notice that both of the models are able to recognize anomalies during the preictal and postictal phases.
Specifically, SincVAE identifies 18.04 % of anomalies in the preictal phase compared to 7.49 % by VAE, and it detects 50 % of anomalies in the postictal phase, against the 27.81 % by VAE. This enhanced capability to detect anomalies during the preictal and postictal phases suggests that SincVAE could be particularly advantageous for applications that require an early warning system for seizures and consistent monitoring during the recovery phase following a seizure.
| Subject | Preictal | Postictal | ||
@@
Table 11 extends the analysis shown in Figure 10 covering all subjects but specifically focusing on the preictal and postictal phases. The anomalies detected during these phases are averaged across the seizure tracks for each subject, since vary in numerosity. The findings from Figure 10 are consistent across the majority of subjects, indicating that SincVAE detects a higher rate of anomalies during the preictal and postictal phases.

### v2

@@ 5 Results
In this section, the results of the seizure detection experiments conducted on both the Bonn and CHB-MIT datasets are presented. For each dataset, the models were trained using the configurations identified during their respective model selection stages. During each training phase, of the training data was randomly sampled and used as the validation set.
It is worth noting that for each result, the model without the SincNet layer is referred to as VAE; whereas, the model incorporating a SincNet layer on top of the encoder network, which is the proposal of this work, is referred to as SincVAE. It is crucial to mention that the primary focus of these experiments is to integrate SincNet into the VAE framework and evaluate its effectiveness in addressing the problem, rather than achieving new state-of-the-art results on the datasets.
@@ 5.1 The Bonn Dataset
The results of the seizure detection for both declared comparisons, i.e. Set A vs Set E and Set B vs Set E, will be discussed in this subsection. In both scenarios, the test set includes 2760 samples, which exhibits a significant imbalance with 460 samples labeled as non-seizure and 2300 labeled as seizure. This imbalance was taken in account during the evaluations.
Figure 8 shows the MSE distributions for the seizure test data (top-left), non-seizure test data (top-right), and the validation data (bottom), comparing the performance of the VAE (orange) and SincVAE (blue) models. Both models produce right-skewed distributions, consistent with the optimization process focused on minimizing reconstruction error. For the non-seizure test data, the models exhibit similar performance, with the VAE showing a slightly lower mode. This behavior aligns with the distribution observed in the validation set, demonstrating the models’ effectiveness in detecting non-seizure patterns. However, on the seizure data, the VAE model demonstrates better reconstruction performance, while the SincVAE model exhibits higher variability, evident through its pronounced right tail and elevated reconstruction errors. SincVAE’s larger discrepancy between its seizure and non-seizure MSE values could indicate a better distinction between the seizure and non-seizure conditions, which could ease the threshold-based classification. In contrast, the narrower MSE range of VAE could make this differentiation more challenging. Similar conclusions can be drawn on Set B (see Figure 9).
To assess the classification performance of the models, a classification threshold must be determined. This threshold is typically based on the MSE values derived from the trained model on the validation or training set, as suggested by various studies [111, 112, 113], and tailored to meet specific user requirements. For this study, the classification threshold was selected using the following criteria:
@@
| Method | Threshold set to (%) | Threshold set to (%) | ||||
| F1 | Precision | Recall | F1 | Precision | Recall | |
| SincVAE | 99.8 | 100 | 99.6 | 99.4 | 98.9 | 100 |
| VAE | 95.6 | 100 | 91.7 | 99.2 | 99.2 | 99.2 |
@@
Considering as true positive an anomalous sample classified as seizure class, and as true negative a non-anomalous sample classified as non-seizure class, the classification performance of the models over the test set was evaluated using the following metrics [70]:
1.
Recall, defined as:
(11) that measures the ability of the model to correctly identify all positive instances (e.g., seizure events);
2.
Precision, defined as:
(12) that measures the accuracy of the positive predictions made by the model;
3.
F1, defined as:
(13) that provides a single metric that captures both the completeness and accuracy of the model’s predictions by balancing the trade-off between recall and precision.
In this work, the F1 score was chosen as the primary metric because it offers a balanced evaluation of the model’s performance by accounting for both recall and precision. This makes it especially appropriate for scenarios involving imbalanced data [70]. Figure 10 shows the confusion matrices and F1, Precision and Recall scores for both SincVAE and VAE, evaluated under the two thresholds and , on the test data for the Set A vs Set E case. By focusing on the confusion matrices, both models demonstrate to correctly identify non-seizure instances, as reflected by high values of true negatives across both thresholds. VAE tends to generate fewer false positives: using the results are comparable, indeed SincVAE detects the of the non-seizure data as seizure against the detected by VAE, while using SincVAE detects the of the non-seizure data as seizure against the detected by VAE. Notably, SincVAE shows fewer false negatives compared to the VAE model: using , SincVAE detects the of the seizure data as non-seizure against the detected by VAE, while using , SincVAE detects the of the seizure data as non-seizure against the detected by VAE. This suggests that SincVAE may be more precise in predicting seizures, aligning with the observations made in Figure 8 and Figure 9. In terms of scores, the performance of SincVAE appears to be superior to that of VAE on this experimental case: SincVAE demonstrates higher F1 scores, indicating it is more reliable for seizure detection across the chosen thresholds. Specifically, SincVAE shows improvements of F1 on and F1 on .
| Method | Threshold set to (%) | Threshold set to (%) | ||||
@@
| SincVAE | 96.7 | 100 | 93.6 | 98.9 | 98.5 | 99.4 |
| VAE | 94.8 | 100 | 90.2 | 98.0 | 98.0 | 97.9 |
@@
Regarding the case Set B vs Set E, Figure 11 shows the confusion matrices and the same classification scores. By focusing on the confusion matrices, SincVAE consistently maintains high performance across both thresholds, demonstrating a lower number of both false positives and false negatives. In terms of false negatives, using , SincVAE detects the of the seizure data as non-seizure, against the detected by VAE; using , SincVAE detects the of the seizure data as non-seizure, against the detected by VAE. In terms of false positives instead, using the performance is comparable, indeed SincVAE detects the of the non-seizure data as seizure against the detected by VAE; using , SincVAE detects the of the non-seizure data as seizure, against the detected by VAE. This indicates a more robust ability of SincVAE to accurately differentiate between seizure and non-seizure events. By focusing on the classification scores, the SincVAE model consistently outperforms in terms of F1 score and recall across both scenarios and thresholds, indicating robust predictive capabilities, particularly in minimizing missed seizures, which could be a critical aspect for clinical applications, due to its impact on enhancing patient safety and treatment efficacy. In particular, SincVAE shows improvements of F1 on and F1 on . Figure 12 shows the receiver operating characteristic (ROC) [70] curve and the area under the curve (AUC) [70] of SincVAE and VAE over the two cases in analysis. The ROC curve is a graphical representation of a classifier’s performance across different threshold values. It plots the True Positive Rate (TPR) against the False Positive Rate (FPR), defined as:
The AUC quantifies the overall performance of the model by measuring the area under the ROC curve. In percentage, an AUC value of indicates a perfect classifier, while a value of suggests no better performance than random chance. Since AUC is threshold-independent, it is particularly useful for comparing models and evaluating their discriminative ability across varying decision boundaries. From Figure 12, it can be noticed that, in the case Set A vs Set E, SincVAE reaches an AUC of while VAE an AUC of , while in the case Set B vs Set E, SincVAE reaches an AUC of while VAE an AUC of . These results confirm that SincVAE consistently outperforms VAE in terms of AUC, highlighting its superior ability to discriminate between the classes in both scenarios.
Figure 14 shows instead the EEG signal reconstructions performed by both the VAE and SincVAE models on seizure samples from Set E. Notably, the VAE model exhibits superior reconstruction fidelity for these seizure signals compared to SincVAE, which results in a higher number of false negatives. Consistent with the classification results, this difference in reconstruction quality suggests varying performance between the two models in processing seizure-related EEG data. The SincVAE’s difficulty in reconstructing seizure signals suggests its ability to distinguish between normal and seizure states, thereby establishing it as a more reliable solution for seizure detection.
Finally, Figure 15 shows the 16 filters learned by the SincNet layer in the SincVAE architecture after the training stage, in the time domain. These filters are visualized only for the network trained on Set A, as the network trained on Set B shows similar characteristics. Details regarding the learned filters, including the low-frequency, high-frequency, and bandwidth, are reported in Table 5. Figure 16 shows the leftmost non-seizure signal shown in Figure 13 decomposed by the SincNet learned filters as an example of the filtering process.
| LF (Hz) | HF (Hz) | Bandwidth (Hz) |
| 0.00 | 3.11 | 3.11 |
| 2.80 | 5.69 | 2.89 |
| 5.08 | 7.71 | 2.62 |
| 7.31 | 9.65 | 2.34 |
| 9.81 | 12.21 | 2.40 |
| 12.44 | 14.94 | 2.50 |
| 14.96 | 17.48 | 2.52 |
| 17.40 | 19.83 | 2.43 |
| 19.68 | 21.94 | 2.25 |
| 22.36 | 24.83 | 2.47 |
| 24.85 | 27.25 | 2.40 |
| 27.38 | 29.83 | 2.46 |
| 29.99 | 32.51 | 2.52 |
| 32.30 | 34.68 | 2.38 |
| 35.18 | 37.96 | 2.78 |
| 37.51 | 40.10 | 2.59 |
5.1.1 Comparison with Existing State-of-the-Art Anomaly Detection Methods
To further assess the effectiveness of our proposed approach, we provide a comparison with several existing state-of-the-art anomaly detection methods. It is important to remark that our goal is not to establish new state-of-the-art results over the dataset under investigation. Instead, we are focused on offering a meaningful comparison with other widely recognized models, giving the reader a clearer idea of how our method performs in relation to existing techniques in the field. Thus, we have chosen to evaluate the performance of several widely used anomaly detection methods to provide a robust comparison of our approach. Specifically, we include one-class support vector machine (OC-SVM) [6], isolation forest (IF) [114], and local outlier factor (LOF) [115]. Additionally, we used PCA [70], where the classification was performed by thresholding on the MSE over the reconstruction, consistent with the approach used for SincVAE in our analysis. In particular, only the threshold was used in this case. Each method was optimized through a model selection stage, following the same details applied on the SincVAE optimization. We employed random search [28] to explore various hyperparameter configurations for each algorithm, conducting 15 trials per method to ensure robust evaluation. The results corresponding to the best model obtained from each algorithm are reported in Table 6.
| Method | Set A vs Set E | Set B vs Set E | ||||
@@ 5.1.1 Comparison with Existing State-of-the-Art Anomaly Detection Methods
| SincVAE | 99.4 | 98.9 | 100 | 98.9 | 98.5 | 99.4 |
| VAE | 99.2 | 99.2 | 99.2 | 98.0 | 98.0 | 97.9 |
| OC-SVM | 98.5 | 97.0 | 100 | 98.4 | 97.2 | 99.6 |
| IF | 99.3 | 99.2 | 99.5 | 95.7 | 98.1 | 93.6 |
| LOF | 99.2 | 98.5 | 100 | 98.7 | 97.6 | 99.8 |
| PCA | 96.9 | 99.2 | 94.7 | 94.6 | 98.0 | 91.4 |
The results show that SincVAE achieves the highest F1 score in both cases, i.e. for Set A vs Set E and for Set B vs Set E. This highlights the robustness of our approach in detecting anomalies while maintaining an optimal balance between Precision and Recall. Additionally, in terms of Recall, SincVAE achieves the best Recall on the Set A vs Set E case, that is , while on the Set B vs Set E case its Recall is from the best, achieved by LOF; in terms of Precision, SincVAE reaches a value of from the best, achieved by VAE, IF and PCA, while on the Set B vs Set E it shows the best value, that is . Given the consistent superior performance of SincVAE, all subsequent experiments presented in this manuscript will be centered solely on the comparison between SincVAE and VAE. This focus allows for a more in-depth evaluation of the advantages offered by SincVAE in seizure detection, while maintaining a clear and direct comparison keeping only VAE results as baseline.
@@ 5.2 The CHB-MIT Dataset
| Subject | F1 (%) | Precision (%) | Recall (%) | ROC-AUC (%) | ||||
| SincVAE | VAE | SincVAE | VAE | SincVAE | VAE | SincVAE | VAE | |
| 1 | 91.9 | 91.8 | 90.0 | 90.9 | 93.9 | 92.8 | 97.5 | 97.2 |
| 2 | 75.6 | 70.8 | 65.1 | 62.7 | 90.1 | 81.3 | 97.2 | 95.8 |
| 3 | 95.0 | 94.3 | 92.7 | 92.2 | 97.5 | 96.5 | 98.5 | 97.9 |
| 4 | 71.8 | 64.5 | 64.4 | 61.7 | 81.2 | 67.5 | 91.7 | 88.9 |
| 5 | 95.6 | 93.4 | 94.6 | 94.2 | 96.6 | 92.7 | 99.2 | 98.4 |
| 6 | 10.8 | 6.9 | 20.8 | 16.7 | 7.2 | 4.3 | 79.7 | 79.3 |
| 7 | 96.2 | 93.1 | 98.1 | 98.0 | 94.5 | 88.6 | 99.7 | 99.6 |
| 8 | 76.8 | 79.9 | 80.8 | 84.7 | 73.1 | 75.6 | 80.4 | 84.5 |
| 9 | 84.9 | 92.3 | 74.7 | 89.6 | 98.4 | 95.2 | 99.1 | 97.6 |
| 10 | 95.9 | 87.8 | 99.8 | 99.4 | 92.4 | 78.5 | 99.8 | 99.5 |
| 11 | 97.5 | 96.8 | 98.6 | 98.7 | 96.4 | 94.9 | 99.3 | 99.1 |
| 12 | 17.5 | 16.1 | 43.8 | 30.4 | 10.9 | 10.9 | 76.1 | 74.5 |
| 13 | 15.7 | 16.5 | 70.4 | 71.4 | 8.8 | 9.3 | 66.9 | 72.0 |
| 14 | 30.8 | 30.8 | 50.9 | 50.9 | 22.0 | 22.0 | 55.9 | 55.3 |
| 15 | 66.0 | 60.6 | 92.6 | 92.2 | 51.3 | 45.2 | 88.0 | 88.8 |
| 16 | 5.5 | 5.3 | 4.2 | 4.0 | 7.9 | 7.9 | 69.5 | 70.7 |
| 17 | 20.7 | 9.6 | 94.4 | 83.3 | 11.6 | 5.1 | 54.1 | 53.9 |
| 18 | 85.3 | 85.4 | 94.6 | 95.0 | 77.6 | 77.6 | 95.4 | 95.3 |
| 19 | 85.9 | 88.1 | 77.9 | 82.8 | 95.8 | 94.1 | 98.4 | 98.2 |
| 20 | 69.8 | 67.0 | 87.2 | 88.4 | 58.2 | 53.9 | 83.4 | 83.6 |
| 21 | 26.6 | 9.1 | 54.7 | 50.0 | 17.6 | 5.0 | 77.7 | 70.9 |
| 22 | 92.7 | 90.2 | 95.3 | 95.1 | 90.2 | 85.8 | 99.2 | 97.9 |
| 23 | 81.7 | 85.4 | 89.5 | 94.6 | 75.2 | 77.9 | 98.6 | 99.0 |
| 24 | 67.0 | 68.3 | 70.8 | 75.2 | 63.5 | 62.6 | 69.8 | 69.6 |
@@
Based on the findings from the experiments with the Bonn dataset, the analyses on the CHB-MIT dataset will be presented in terms of F1, Precision, Recall and ROC-AUC performances of both the SincVAE and VAE models for each subject in the CHB-MIT dataset. The decision threshold was chosen to be unique and equal to the 95th percentile of the validation set’s MSE for each subject. Results are shown in Table 7.
@@
Observing the F1 scores, SincVAE performs comparably or slightly better than VAE for the majority of the subjects in the CHB-MIT dataset. Specifically, for subjects 1, 2, 3, 4, 5, 6, 7, 10, 11, 12, 15, 16, 17, 20, 21, and 22, SincVAE either matches or exceeds the performance of VAE. Notably, there are cases where VAE has a higher F1 score than SincVAE, specifically for subjects 8, 9, 13, 18, 19, 23, and 24. These variations suggest that VAE might be better suited for certain cases. It is also important to note that, according to the ROC-AUC results, SincVAE outperforms VAE in 18 out of 24 cases, further validating its effectiveness over VAE.
@@
It is worth noting that the F1 scores for some subjects, such as 6 and 13, are low, which may be influenced by factors such as the inherent quality of the EEG recordings, the need for specific data preprocessing or the need for more sophisticated model architectures. However, the main aim of this study is to explore the effectiveness of integrating a SincNet layer within a VAE framework, rather than establishing new state-of-the-art results on the CHB-MIT dataset. Thus, discussions on specific architectural enhancements to improve EEG data quality are outside the scope of this work, which focuses instead on assessing the added value of the SincNet layer within the VAE architecture.
Finally, to assess whether SincVAE outperforms VAE from a statistical significance perspective, hypothesis testing was conducted, first considering F1 scores and then ROC-AUC values, with the significance level set at . For the F1 scores, the null hypothesis stated that the mean difference between the performance of SincVAE and VAE was zero, while the alternative hypothesis suggested that SincVAE outperforms VAE, i.e., the mean difference is greater than zero. To check for normality in the differences, the Shapiro-Wilk test was first applied. The p-value obtained was , indicating that the differences followed a normal distribution. Consequently, a one-sided paired t-test [116] was performed to test whether the mean difference between SincVAE and VAE was greater than zero. The resulting p-value, , was below the significance level , leading to the rejection of the null hypothesis and the conclusion that SincVAE significantly outperforms VAE. For the ROC-AUC values, a similar approach was followed. The Shapiro-Wilk test yielded a p-value of , indicating that the differences did not follow a normal distribution. Therefore, the one-sided Wilcoxon signed-rank test [116] was applied. The p-value from the Wilcoxon test was , which is below the significance level . As a result, the null hypothesis was rejected, and it was concluded that SincVAE significantly outperforms VAE in terms of ROC-AUC values.
@@
As stated above, the seizure tracks in the CHB-MIT dataset include preictal, ictal, and postictal phases. Figure 17 shows the MSE values obtained for each second of track number 19 from subject 9, a track identified by the dataset’s authors as including a seizure event from seconds 5299 to 5361. These specific seconds are visually demarcated with two dashed vertical blue lines. The horizontal red line on the plot indicates the decision threshold, set at the 95th percentile of the MSE values obtained from the validation set. Blue dots on the plot represent EEG recordings that have been classified as non-seizure, while red dots indicate those classified as seizure.
| Subject | Preictal (%) | Postictal (%) | ||
@@
It can be noted that several recordings are classified as anomalous by both of the models either in the preictal or in the postictal phase. Figure 18 shows the percentage of anomalous points detected by SincVAE and VAE in the preictal, ictal and postictal phases. Both models detect seizures during the ictal phase with a high rate, and demonstrate effectiveness in recognizing the interictal phase with a low rate of false positives. It is interesting to notice that both of the models are able to recognize anomalies during the preictal and postictal phases. Specifically, SincVAE identifies of anomalies in the preictal phase compared to by VAE, and it detects of anomalies in the postictal phase, against the by VAE. This enhanced capability to detect anomalies during the preictal and postictal phases suggests that SincVAE could be particularly advantageous for applications that require an early warning system for seizures and consistent monitoring during the recovery phase following a seizure. Table 8 extends the analysis shown in Figure 18 covering all subjects but specifically focusing on the preictal and postictal phases. The anomalies detected during these phases are averaged across the seizure tracks for each subject, as their numerosity varies. The findings from Figure 18 are consistent across the majority of subjects, indicating that SincVAE detects a higher rate of anomalies during the preictal and postictal phases.